Sam lives in Houston, TX and Jack lives in Denver, Co. Use the CPI for Houston and Denver areas in Table 1.1. below to fill in the blank cells in Table 1.2.  Show your work by typing in the formulas you use.

Table 1.1. Consumer Price Index (CPI-U) All items

Base Period:   2009=100

Year

Houston-the Woodlands-Sugar Land, TX

Denver-Aurora-Lakewood, CO

2009

100.00

100.00

2010

101.93

101.87

2011

105.25

105.63

2012

107.20

107.68

2013

108.97

110.67

2014

112.01

113.74

2015

111.83

115.08

2016

113.61

118.27

2017

115.83

122.27

2018

118.60

125.61

Table 1.2. Earnings and CPI Annual Averages, Houston, TX and Denver, CO

2017

2018

Sam earned in Houston, TX (current dollars)

$55,000

$59,000

Jack earned in Denver, CO (current dollars)

$57,000

$60,000

CPI-U All Items for Houston

CPI-U All Items for Denver

Sam's real earnings (inflation-adjusted) in Houston, expressed in 2009 dollars

Jack’s real earnings (inflation-adjusted) in Denver, expressed in 2009 dollars
Answer #1

In case of Sam,

Real Earnings in 2017=Earning in 2017(Current dollars)*(CPI in base 2009/CPI in 2017)

=55000*(100/115.83)=$47483.38

Real Earnings in 2018=Earning in 2018(Current dollars)*(CPI in base 2009/CPI in 2018)

=59000*(100/118.60)=$49747.05

In case of Jack,

Real Earnings in 2017=Earning in 2017(Current dollars)*(CPI in base 2009/CPI in 2017)

=57000*(100/122.27)=$46618.14

Real Earnings in 2018=Earning in 2018(Current dollars)*(CPI in base 2009/CPI in 2018)

=60000*(100/125.61)=$47766.90

2017 2018
Sam earned in Houston, TX (current dollars) 55000 59000
Jack earned in Denver, CO (current dollars) 57000 60000
CPI-U All Items for Houston 115.83 118.6
CPI-U All Items for Denver 122.27 125.61
Sam's real earnings (inflation-adjusted) in Houston, expressed in 2009 dollars 47483.38 49747.05
Jack’s real earnings (inflation-adjusted) in Denver, expressed in 2009 dollars 46618.14 47766.90